OpenStudy (anonymous):

Choose the correct simplification of the expression (3xy4)2(y2)3. 6x2y14 9x2y14 9x3y11 6x3y11

OpenStudy (anonymous):

....? lol

OpenStudy (anonymous):

G2g so sry can't help right now :(

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

\[(3xy^4)^{2} \ (y^2)^{3}\]

OpenStudy (anonymous):

(a^b)^c=a^bc

OpenStudy (zepp):

Expand \((3xy^4)^2 into (3*x*y^2*3*x*y^2)\)

OpenStudy (zepp):

And you'll get \(\large 3^2*x^2*(y^2)^2*(y^2)^3\)

OpenStudy (campbell_st):

use the index law for power of a power \[(x^a)^b = x^{a \times b}\] so you have \[3^2 x^{1\times2}y^{4\times2} \times y^{2\times3}\] simplify the powers and evaluate 3^2 last set will be same base when multiplying \[x^a \times x^b = x^{a+b}\] hope this helps

OpenStudy (anonymous):

9x^2y^14 is the correct simplification

OpenStudy (anonymous):

thank you ^^

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

teehee there you go :)